OpenStudy (anonymous):

The container that holds the water for the football team is 1/3 full. After pouring out 6 gallons of water, it is 1/9 full. How many gallons can the container hold?

OpenStudy (phi):

think like this: let "x" be the number of gallons when full. How many gallons in the container when 1/3 full? you write x/3 (that means we divide the number of gallons by 3)

OpenStudy (anonymous):

alright..

OpenStudy (phi):

now pour out 6 that is like "subtracting" x/3 - 6

OpenStudy (phi):

what we now have is it is 1/9 full. any idea how to write that using x?

OpenStudy (anonymous):

=x/9?

OpenStudy (phi):

yes. we have \[ \frac{x}{3} - 6 = \frac{x}{9} \] now we switch from "what is going on mode" to "algebra mode" to solve I would multiply both sides (and all terms) by 9 (this will get rid of the fractions)

OpenStudy (anonymous):

if i do x/3-6=x/9 then x would equal 27

OpenStudy (phi):

the point of doing this is not the answer, it is how you get the answer.

OpenStudy (anonymous):

okay so I would multiply each side by 9 and get rid of things then it would be 3 multiplied by 9

OpenStudy (phi):

like this \[ 9 \cdot \frac{x}{3} - 9 \cdot 6 = 9 \frac{x}{9} \]

OpenStudy (anonymous):

so then because it equals 27 (3 times 9 does) would i subtract 6 and have the answer of 21?

OpenStudy (phi):

the first term \[ 9\cdot \frac{x}{3} \] or \[ \frac{9 \cdot x}{3} \] you can divide 3 into 9

OpenStudy (phi):

or, 9/3 is ?

OpenStudy (anonymous):

wouldnt it be 9 times x/3

OpenStudy (phi):

yes 9 times x/3 which you write as \[ 9 \cdot \frac{x}{3} \] or \[ \frac{9\cdot x}{3} \] or \[ \frac{9}{3} \cdot x \] all different ways of writing the same thing. but the important part is it means you can divide 3 into 9

OpenStudy (phi):

Is this confusing? If it is , it is worth figuring it out, because algebra uses this a *lot*

OpenStudy (anonymous):

yes im super confused

OpenStudy (phi):

If you have time, you can learn the idea You know how to figure out \[ \frac{4}{2} \] = 2 right?

OpenStudy (phi):

4 is the same as 2*2 so we could write the problem as \[ \frac{2\cdot 2}{2} \] and we know the answer is still 2

OpenStudy (phi):

or , another example, \[ \frac{30}{5} = \frac{2\cdot 3\cdot 5}{5} \] if you divide 5/5 you get \[ \frac{2\cdot 3\cdot \cancel{5}}{\cancel{5}} = \frac{2\cdot 3}{1}= 6\]

OpenStudy (phi):

and you know 30/5 = 6 we got the correct answer. we use that same trick with \[ \frac{9 x}{3} = \frac{3 \cdot 3 \cdot x}{3} \] or, using the trick \[ \frac{\cancel{3} \cdot 3 \cdot x}{\cancel{3}} = \frac{3x}{1} = 3x\]